OpenStudy (jiteshmeghwal9):

Find the amount and interest on Rs.24000 for 9 months, being given that interest is compounded quarterly at the rate of 20 paisa a rupee per month.

OpenStudy (jiteshmeghwal9):

@ganeshie8 sir plz help :)

ganeshie8 (ganeshie8):

we can use this formula whenever we need to work out 'compounded' interest : \(A = P(1+\frac{r}{n})^{nt} \) A = amount of money accumulated after n years, including interest. P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. n = number of times the interest is compounded per year

OpenStudy (jiteshmeghwal9):

sir can u give me long step method it's compulsory :)

ganeshie8 (ganeshie8):

ha

ganeshie8 (ganeshie8):

like the one we worked out earlier problem ?

OpenStudy (jiteshmeghwal9):

no, i mean finding C.I through Simple interest of one one year

ganeshie8 (ganeshie8):

..... rate of 20 paisa a rupee per month. here we're given SI of one month, is that right

OpenStudy (jiteshmeghwal9):

like if we r said to find the C.I of 3 years at the rate of 8%p.a. then finding the S.I. o 1 yr & then find amount

OpenStudy (jiteshmeghwal9):

i don't know :(

ganeshie8 (ganeshie8):

i get u, here its not for one year thats given, its for 1 month as it says '20 paisa a rupee per month'

OpenStudy (jiteshmeghwal9):

yeah !

ganeshie8 (ganeshie8):

since there are 12 months in a year, the rate per year would be ?

OpenStudy (jiteshmeghwal9):

1.20 times 12 ?

ganeshie8 (ganeshie8):

20 paisa a rupee per month = 12 x 20 paisa a rupee per year = 3 x 20 paisa a rupee per quarter.

OpenStudy (jiteshmeghwal9):

ya !

ganeshie8 (ganeshie8):

all 3 are same ok, we will take the last one as we are dealing with quarters

OpenStudy (jiteshmeghwal9):

k!

ganeshie8 (ganeshie8):

every quarter, 3 x 20 = 60 paisa for every rupee is added as interest. which is nothing but 60% rate per quarter

ganeshie8 (ganeshie8):

60 paisa per 1 rupee = 60%

ganeshie8 (ganeshie8):

make sure its clear :)

OpenStudy (jiteshmeghwal9):

i don't gt it surely

ganeshie8 (ganeshie8):

every quarter, 3 x 20 = 60 paisa for every rupee is added as interest.

ganeshie8 (ganeshie8):

that statement is clear right ?

OpenStudy (jiteshmeghwal9):

ya !

ganeshie8 (ganeshie8):

60 paisa for every rupee is what percent ?

ganeshie8 (ganeshie8):

us this : 1 rupee = 100 paisa

OpenStudy (jiteshmeghwal9):

60/100 times 100

ganeshie8 (ganeshie8):

thats right !!

OpenStudy (jiteshmeghwal9):

60 %

ganeshie8 (ganeshie8):

so, 60% is our quarterly rate.

ganeshie8 (ganeshie8):

lets work out the rest now

OpenStudy (jiteshmeghwal9):

ok !

ganeshie8 (ganeshie8):

Rs.24000 for 9 months

ganeshie8 (ganeshie8):

9 months = 3 quarters => they're calculating interest 3 times.

OpenStudy (jiteshmeghwal9):

yes !

ganeshie8 (ganeshie8):

total amount with interest = \(A(1+r)^n\) r = 60% = 60/100 = 6/10 = 3/5 n = 3

ganeshie8 (ganeshie8):

\(24,000(1+\frac{3}{5})^3\) = ?

OpenStudy (jiteshmeghwal9):

\[24000\left( \frac{8 \times 8 \times 8}{5 \times 5 \times 5} \right)\]\[{24000 \times 512 \over125}=98304\]

ganeshie8 (ganeshie8):

perfect ! so that the amount including interest.

OpenStudy (jiteshmeghwal9):

yes !

ganeshie8 (ganeshie8):

find out the interest component by subtracting the initial money 24,000

OpenStudy (jiteshmeghwal9):

74,304

ganeshie8 (ganeshie8):

so whats the lesson.... ? compound interest is really very very nasty ! we started with 24,000 and in just 9 months, we are forced to pay interest of 74,304 !!!